OpenStudy (anonymous):

Solve the system using any method. -4x-5y-z=18 -2x-5y-2z=12 2x+5y+2z=4

OpenStudy (amistre64):

-2x-5y-2z=12 2x+5y+2z=4 -------------- 0+0+0 = 16 is false so there isnt a solution that i can see

OpenStudy (amistre64):

you have 2 parallel lines and a cross between them; there is no solution

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

 something like this :)

OpenStudy (anonymous):

thnaks so much